The smallest angle of a triangle measures 44 degrees less than the largest angle. The sum of the two smaller angles is 20 degrees more than the measure of the largest angle. Find the measures of all three angles.
Answer by shree840(260)

largest angle x
smallest angle x-44
middle +smallest=x+20
x+20+x=180
2x=160
x=80
smallest =80-44=36
middle 100-36=64 ANS